# Navigating the Reality Gap: Privacy-Preserving Adaptation of ASR for Challenging Low-Resource Domains

### v2

@@ Abstract
Automatic Speech Recognition (ASR) holds immense potential to assist in clinical documentation and patient report generation, particularly in resource-constrained regions. However, deployment is currently hindered by a technical deadlock: a severe “Reality Gap” between laboratory performance and noisy, real-world clinical audio, coupled with strict privacy and resource constraints. We quantify this gap, showing that a robust multilingual model (IndicWav2Vec) degrades to a 40.94% WER on rural clinical data from India, rendering it unusable. To address this, we explore a zero-data-exfiltration framework enabling localized, continual adaptation via Low-Rank Adaptation (LoRA). We conduct a rigorous investigative study of continual learning strategies, characterizing the trade-offs between data-driven and parameter-driven stability. Our results demonstrate that multi-domain Experience Replay (ER) yields the primary performance gains, achieving a 17.1% relative improvement in target WER and reducing catastrophic forgetting by 55% compared to naive adaptation. Furthermore, we observed that standard Elastic Weight Consolidation (EWC) faced numerical stability challenges when applied to LoRA in noisy environments. Our experiments show that a stabilized, linearized formulation effectively controls gradient magnitudes and enables stable convergence. Finally, we verify via a domain-specific spot check that acoustic adaptation is a fundamental prerequisite for usability which cannot be bypassed by language models alone.
Navigating the Reality Gap: Privacy-Preserving Adaptation of ASR for Challenging Low-Resource Domains
@@ 1 Introduction
The recent surge in Self-Supervised Learning (SSL) has propelled Automatic Speech Recognition (ASR) to near-human performance on standardized benchmarks. Foundational models like Meta’s Wav2Vec 2.0 baevski2020wav2vec and OpenAI’s Whisper radford2023robust promise a future where automated transcription can digitize patient reports, allowing clinics to improve throughput with reduced operational costs. However, for specialized, high-impact domains such as rural healthcare and telemedicine, this promise remains unfulfilled. The "reality gap," which represents the disparity between clean training corpora and the chaotic, noisy, and privacy-constrained environments of real-world clinics, renders these state-of-the-art models practically unusable.
@@
Our baseline analysis reveals that even robust multilingual models like IndicWav2Vec javed2022indicwav2vec degrade to a prohibitively high 40.94% Word Error Rate (WER) when exposed to real-world clinical audio from rural India gram_vaani. This failure is compounded by a technical deadlock: patient data privacy laws preclude the use of cloud-based adaptation services nawaz2019privacy, while rural infrastructure lacks the high-end compute required for traditional model retraining. This creates a scenario where models cannot improve because data cannot leave the local environment, and localized compute is too constrained for standard fine-tuning.
@@
To break this deadlock, we propose a privacy-preserving adaptation framework for localized, stream-based learning. We define "privacy-preserving" strictly as data residency: ensuring that raw patient audio never leaves the local device. By leveraging Low-Rank Adaptation (LoRA) hu2022lora, we enable the model to fine-tune on incoming data streams using a fraction of the trainable parameters. To isolate the effects of acoustic adaptation and ensure feasibility on extreme edge devices (e.g., 35W mobile GPUs), we deliberately exclude external Language Models, allowing us to quantify the upper bound of purely acoustic plasticity in this domain. However, sequential adaptation introduces the risk of catastrophic forgetting, where the model loses its foundational linguistic capabilities mccloskey1989catastrophic. To counteract this, we integrate a multi-domain experience replay mechanism chaudhry2019tiny, interleaving small buffers of general-domain data with the incoming clinical stream to anchor the learning trajectory.
@@
Furthermore, we document the challenges encountered when incorporating parameter-regularization (EWC) into this pipeline. We find that standard quadratic formulations can lead to numerical instability and gradient explosion in noisy clinical environments. We empirically demonstrate that a stabilized, linearized importance estimation (Absolute Fisher) hsu2023absolute aids robust convergence in LoRA-based continual adaptation. Our results demonstrate that these strategies yield a 17.1% relative reduction in WER on the target domain, effectively stabilizing the model’s performance in a privacy-constrained setting. The primary research contributions of our work are:
  • Quantification of the Reality Gap: We establish the 40.94% WER floor for state-of-the-art multilingual models in rural clinical environments, providing a rigorous benchmark for domain adaptation.
  • Privacy-Preserving Adaptation Framework: We develop a LoRA-based on-device adaptation pipeline that enables continuous learning from data streams while ensuring 100% patient data residency and computational efficiency.
  • Investigation of Architectural Stability: We diagnose the stability challenges of standard EWC when integrated with LoRA for noisy domain adaptation and empirically validate the benefits of linearized Fisher importance for long-term pipeline health.
  • Empirical Validation on Low-Resource Data: We demonstrate a 17.1% relative improvement in WER on the Gram Vaani dataset, establishing a viable pathway for self-improving ASR systems in high-impact environments.
@@ 2.1 Self-Supervised Acoustic Foundations and the Domain Gap
Despite these capabilities, a “usability gap” exists between laboratory benchmarks and real-world deployment. radford2023robust demonstrate that foundational models suffer severe degradation on low-resource languages, rendering raw transcripts unusable. This is exacerbated in rural healthcare by noisy telephonic audio and diverse regional dialects (gram_vaani). Standard pre-trained models struggle to generalize to these conditions without targeted adaptation. Furthermore, privacy constraints preclude cloud-based services (nawaz2019privacy), creating a deadlock where models cannot improve because data cannot leave the local environment.
@@ 2.2 Parameter-Efficient Fine-Tuning (PEFT) and the Efficiency Gap
Bridging the domain gap typically requires fine-tuning on target data. However, for Large Audio Models (LAMs), full fine-tuning is computationally prohibitive and prone to overfitting, particularly with scarce data. This presents an efficiency gap: the high-end GPUs required for full fine-tuning are unavailable on the resource-constrained edge devices of rural hospitals.
[…]
2.3 Continual Learning and the Stability Gap
While PEFT solves efficiency, self-improving systems face the stability gap of Continual Learning (CL), specifically Catastrophic Forgetting, where new training erodes previously learned capabilities (mccloskey1989catastrophic; parisi2019continual). In ASR, naive updates on recent transcripts lead to overfitting on specific speakers or acoustic conditions.
To mitigate this, we employ regularization methods rooted in Elastic Weight Consolidation (EWC) (kirkpatrick2017overcoming). While schwarz2018progress proposed Online EWC, standard implementations require computing the Fisher Information Matrix (FIM) using squared gradients, which can lead to parameter “freezing” or divergence in noisy domains (chaudhry2018riemannian). To align with our efficiency and stability constraints, we implement LoRA-EWC (wang2023language), applying regularization specifically to the low-rank adapters. Departing from the standard quadratic FIM, we utilize a Linearized Fisher importance $F_{i}$ based on the absolute value of the gradients ($|g|$), following the “Absolute Fisher” formulation (hsu2023absolute). This modification, combined with a quadratic parameter shift penalty, prevents the rigid locking associated with standard EWC, ensuring the model retains the plasticity needed to adapt to clinical domain shifts without losing foundational knowledge.
Complementing this regularization, we employ Experience Replay (ER), which interleaves a buffer of prior examples with current training data (chaudhry2019efficient). yang2022online demonstrated that replay-based methods perform robustly for ASR adaptation. We utilize a multi-domain experience replay mechanism, blending “clean” general-domain speech with “noisy” domain-specific samples via a prioritized buffer management strategy (Algorithm A.2). This combined strategy ensures the system specializes in the clinical environment while preserving the linguistic competence required for reliable long-term operation.
@@ 3 Methodology
We propose a privacy-preserving, adaptive ASR framework designed to bridge the "reality gap" between general-purpose models and the acoustic constraints of rural healthcare. Our approach focuses on Continual Learning (CL), enabling a pre-trained model to adapt sequentially to incoming clinical data streams ($\mathcal{D}_{stream}$) without data leakage or catastrophic forgetting.
@@ 3.1 Base Acoustic Backbone
To enable efficient adaptation on edge devices, we freeze the base model and inject trainable Low-Rank Adaptation (LoRA) matrices into the query and value projection layers.
@@ 3.2 Data-Driven Stability: Multi-Domain Experience Replay
  • General Domain Anchor ($\mathcal{B}_{gen}$): A gender-balanced subset of high-resource, standard Hindi samples (sourced from Kathbath) to preserve foundational linguistic competence.
@@ 3.3 Parameter-Driven Stability: Elastic Weight Consolidation (EWC)
We compute the importance of each LoRA parameter $\theta_{i}$ using the diagonal of the Fisher Information Matrix ($F$). To ensure robust estimation on small batches, we approximate the Fisher Information using the accumulated absolute gradients rather than squared gradients, inspired by Memory Aware Synapses (MAS) (aljundi2018memory) and the Absolute Fisher metric (hsu2023absolute):
@@ 4.1.1 Metrics
where $S$, $D$, and $I$ represent substitutions, deletions, and insertions, and $N$ is the total count in the reference.
@@ 4.1.2 Datasets
  • Gram Vaani (Target Domain) (gram_vaani): This dataset consists of rural telephonic speech (originally 8kHz, upsampled to 16kHz) and serves as a proxy for the challenging, domain-specific audio encountered in rural hospitals. The content includes medical and agricultural discussions, making it highly relevant for simulating real-world deployment in our target sectors. To strictly simulate a continual learning scenario, we partition the 103 hours of training data into sequential segments, processing them one by one to mimic a live data stream.
@@ 5.2 Stability-Plasticity Pareto Efficiency
  • Data Anchors (ER): Replaying clean source-domain samples (Kathbath) provides an explicit gradient signal that steers the optimization toward a generalized linguistic center. This mechanism proved the most efficient, as evidenced by the V3.1 trajectory on the Pareto frontier.
  • Parameter Anchors (EWC): Implicitly constraining weight drift via EWC protects the model but is “acoustically blind.” It penalizes movement relative to the distribution of the source domain importance without providing the model with the actual acoustic features to reconcile. The superiority of data-driven anchors is further underscored by comparing Single-Domain ER (V2.1) with pure EWC (V4.5). Despite achieving similar target-domain performance, the data-driven rehearsal in V2.1 provided superior general-domain stabilization (+3.33% forgetting) compared to the weight-regularization in V4.5 (+3.58%), confirming that explicit data rehearsal is a more effective stabilizer for low-rank adaptation than implicit parameter constraints.
  • The LoRA-Regularization Bottleneck: We observe a conflict when combining LoRA’s low-rank space with EWC’s parameter-level penalties. In the Hybrid model (V5.1), the cumulative constraints (low rank + EWC penalty) overly restrict the available optimization paths, leading to higher target WER.
5.4 Convergence Dynamics and The “Reality Gap” Floor
A consistent contrast is observed in the convergence behavior across paradigms. While the Naive baseline (V1.1) exhibits erratic, high-variance loss curves, the ER and Hybrid paradigms demonstrate a controlled, monotonic downward trend (Figure 2). This behavior is a critical indicator of long-term viability: it suggests that the system can safely continue to learn from an indefinite data stream without the risk of sudden divergence.
Finally, we observe a consistent convergence around 34% WER given the current data volume. While this suggests that strong strategies like V3.1 could continue improving with additional streaming data, the progression (Figure 2) indicates diminishing returns. This points to a “soft floor” where further gains become increasingly expensive, likely due to the fundamental “Reality Gap”: the acoustic mismatch between 8kHz telephonic recordings and the pre-trained model’s expected 16kHz clarity. This underscores that while localized adaptation can recover significant linguistic performance, overcoming the final acoustic barrier may require architectural enhancements or speech-enhancement front-ends.
@@ 5.5 The Necessity of Acoustic Adaptation
To verify that acoustic adaptation is a prerequisite for usability, we paired both the baseline and adapted models with a domain-specific 4-gram LM trained on Gram Vaani transcripts (Table 2). The unadapted baseline+LM achieves only 34.96% WER, while our V3.1+LM achieves 30.26% WER, a 4.7% absolute improvement. This confirms that LMs alone cannot bridge the reality gap; acoustic adaptation is a fundamental prerequisite for deployment.

### v3

# Navigating the Reality Gap: Privacy-Preserving On-Device Continual Adaptation of ASR for Clinical Telephony

@@ Abstract
Automatic Speech Recognition (ASR) holds immense potential to assist in clinical documentation and patient report generation, particularly in resource-constrained regions. However, deployment is currently hindered by a technical deadlock: a severe “Reality Gap” between laboratory performance and noisy, real-world clinical audio, coupled with strict privacy and resource constraints. Such adaptation is essential for clinical telephony systems, where patient speech is highly variable and transcription errors can directly impact downstream clinical workflows. We quantify this gap, showing that a robust multilingual model (IndicWav2Vec) degrades up to a 40.94% WER on rural clinical telephony speech from India, rendering it unusable. We demonstrate consistent improvements on these helpline interactions without transmitting raw patient data off-device via an on-device continual adaptation framework using Low-Rank Adaptation (LoRA). We conduct an investigative study of stabilization strategies, characterizing the trade-offs between data-driven and parameter-driven approaches. Our results demonstrate that multi-domain Experience Replay (ER) yields the primary performance gains, achieving a 17.1% relative improvement in target WER and reducing catastrophic forgetting by 55% compared to naive adaptation. Furthermore, we investigate a stabilized importance estimation strategy (Absolute Fisher) to ensure robust convergence against the high-variance gradients common in clinical telephony speech. Finally, we verify via a domain-specific spot check that acoustic adaptation is a fundamental prerequisite for usability in healthcare settings which cannot be bypassed by language models alone.
Navigating the Reality Gap: Privacy-Preserving On-Device Continual Adaptation of ASR for Clinical Telephony
@@ 1 Introduction
The recent surge in Self-Supervised Learning (SSL) has propelled Automatic Speech Recognition (ASR) to near-human performance on standardized benchmarks. Foundational models like Meta’s Wav2Vec 2.0 baevski2020wav2vec and OpenAI’s Whisper radford2023robust promise a future where automated transcription can digitize patient reports, allowing clinics to improve throughput with reduced operational costs. However, for specialized, high-impact clinical domains such as rural health helplines and remote clinical services, this promise remains unfulfilled. The "reality gap," representing the disparity between clean training corpora and the chaotic, noisy, and privacy-constrained environments of real-world clinical telephony, renders these state-of-the-art models practically unusable.
@@
Our baseline analysis reveals that even robust multilingual models like IndicWav2Vec javed2022indicwav2vec degrade to a prohibitively high 40.94% Word Error Rate (WER) when exposed to real-world clinical telephony audio from rural India gram_vaani. This failure is compounded by a technical deadlock: patient data privacy laws preclude the use of cloud-based adaptation services nawaz2019privacy, while rural infrastructure lacks the high-end compute required for traditional model retraining. This creates a scenario where models cannot improve because data cannot leave the local environment, and localized compute is too constrained for standard fine-tuning.
@@
To break this deadlock, we propose an on-device adaptation framework for localized, stream-based learning. We focus on data residency: ensuring that raw patient audio never leaves the local device. By leveraging Low-Rank Adaptation (LoRA) hu2022lora, we enable the model to fine-tune on incoming clinical data streams using a fraction of the trainable parameters. However, sequential adaptation in long-running clinical deployments introduces the risk of catastrophic forgetting, where the model loses its foundational linguistic capabilities mccloskey1989catastrophic. While recent benchmarks on real-world Indian speech, such as NIRANTAR javed2025nirantar, indicate that standard regularization methods (EWC, MAS) often fail to prevent this forgetting in diverse domains, we hypothesize that this instability is exacerbated by the acoustic mismatch inherent in low-bandwidth clinical telephony. To counteract this, we integrate a multi-domain experience replay mechanism chaudhry2019tiny, interleaving small buffers of general-domain data with the incoming clinical stream to anchor the learning trajectory.
@@
Furthermore, we investigate parameter-regularization strategies to stabilize this pipeline. While standard quadratic formulations (EWC) can be sensitive to the high-variance gradients of noisy telephony speech, we observe that a balanced importance estimation (Absolute Fisher) hsu2023absolute provides a more resilient regularization signal for this localized application. Our results demonstrate that these strategies yield a 17.1% relative reduction in WER on the target clinical domain, effectively stabilizing the model’s performance while respecting rigid data residency constraints. The primary research contributions of our work are:
  • On-Device Clinical Adaptation: A localized continual adaptation framework for clinical telephony ASR that respects healthcare data residency constraints while maintaining computational efficiency.
  • Empirical Validation on Clinical Speech: Parameter-efficient adaptation (LoRA) combined with multi-domain memory (Replay) yields a 17.1% relative improvement in WER on real-world rural clinical speech without centralized retraining or raw data exfiltration.
  • Investigation of Optimization Robustness: The interplay between data-driven replay and parameter-driven regularization, validating that linearized Fisher importance promotes more reliable convergence in noisy, multi-speaker clinical deployments.
  • Quantification of the Clinical Reality Gap: Establish a rigorous benchmark for 8kHz clinical telephony, providing insights into the technical prerequisites to close the gap between foundational ASR models and frontline healthcare needs.
@@ 2.1 Self-Supervised Acoustic Foundations and the Domain Gap
Despite these capabilities, a “usability gap” exists between laboratory benchmarks and real-world deployment. Clinical ASR errors are particularly risky because they affect medically salient information such as dosages, negations, and symptom reporting finley2018automatic; shafran2003challenges. This is exacerbated in rural healthcare by noisy telephonic audio and diverse regional dialects. Recent large-scale efforts such as NIRANTAR javed2025nirantar have begun characterizing these challenges across 22 Indian languages. We utilize the Gram Vaani ASR dataset (bhanushali2022gramvaani) to represent these challenges; bhanushali2022gramvaani highlight that standard models struggle to generalize to such 8kHz telephony audio without targeted adaptation.
While signal-level enhancement strategies, such as bandwidth expansion and speech super-resolution (lin2023noise; li2019audio), offer an alternative by reconstructing high-frequency components, they introduce significant computational overhead. Deploying a separate BWE model (often GAN or diffusion-based) alongside the ASR system contradicts the strict low-latency requirements of edge devices. Furthermore, privacy constraints preclude cloud-based services (nawaz2019privacy), creating a deadlock where models cannot improve because data cannot leave the local environment.
@@ 2.2 Parameter-Efficient Fine-Tuning (PEFT) and the Efficiency Gap
Bridging the domain gap typically requires fine-tuning on target data. However, for Large Audio Models (LAMs), full fine-tuning is computationally prohibitive and prone to overfitting. This presents an efficiency gap: high-end GPUs are unavailable on the resource-constrained edge devices of rural hospitals.
To address this, we adopt Low-Rank Adaptation (LoRA) (hu2022lora), which enables local fine-tuning on modest clinical hardware. This approach is essential for meeting healthcare Data Residency requirements, as it allows for localized, high-performance adaptation without needing the centralized data storage often required by federated or cloud-based approaches rieke2020future. Recent work by song2024lora on LoRA-Whisper validates this approach for multilingual ASR, demonstrating that PEFT can effectively adapt large Transformers without catastrophic interference. By constraining optimization to a low-dimensional subspace, LoRA resolves deployment challenges in privacy-sensitive clinics.
While PEFT solves efficiency, self-improving systems face the stability gap of Continual Learning (CL), specifically Catastrophic Forgetting, where new training erodes previously learned capabilities (mccloskey1989catastrophic). In healthcare, ensuring model stability over time is a critical safety requirement for clinical evaluation cycles delange2021continual. In ASR, naive updates on recent telephony transcripts lead to overfitting on specific speakers or acoustic conditions. Benchmarks like NIRANTAR javed2025nirantar further underscore this, demonstrating that no single CL strategy currently yields consistent performance across the diverse shifts encountered in real-world Indian speech.
To mitigate this, we employ a hybrid continual learning strategy that synergizes data-level and parameter-level consolidation. First, to address the distribution shift between segments, we implement Experience Replay (ER) (chaudhry2019efficient). Building on recent regularized CL studies for Indic ASR (adethya2025study) and episodic memory frameworks (yang2022online), we utilize a multi-domain experience replay mechanism. While yang2022online focused on gradient projection, we find that a direct data mixing strategy effectively mitigates catastrophic forgetting when combined with PEFT. Our buffer retains a mixture of “hard” examples (high-loss) from the target domain (Gram Vaani) and gender-balanced samples from a clean auxiliary domain (Kathbath).
Complementing this, we constrain parameter drift using Online LoRA-EWC (wang2023language), utilizing the Absolute Fisher importance estimation (hsu2023absolute). We estimate importance using the absolute value of accumulated gradients ($|g|$) rather than the standard squared Fisher Information. While this linear scaling draws inspiration from the numerical stability of Memory Aware Synapses (MAS) (aljundi2018memory), Absolute Fisher explicitly ties the importance to the likelihood of the adaptation task rather than unsupervised output sensitivity. By combining this regularized signal with multi-domain replay, our approach prevents catastrophic forgetting through a dual mechanism: replay maintains the data distribution, while the regularization penalty preserves critical adapter parameters.
@@ 3 Methodology
We propose an on-device, adaptive ASR framework designed to bridge the "reality gap" and ensure deployment-ready reliability within the acoustic constraints of rural clinical telephony. Our approach focuses on Continual Learning (CL), enabling a pre-trained model to adapt sequentially to incoming clinical data streams ($\mathcal{D}_{stream}$) while maintaining strict data residency and preventing catastrophic forgetting.
@@ 3.1 Base Acoustic Backbone
To enable efficient adaptation on edge devices, we freeze the base model and inject trainable Low-Rank Adaptation (LoRA) matrices into the query and value projection layers. This parameter-efficient approach serves as a stability safeguard, preventing large-scale corruption of pre-trained weights while allowing for specialized clinical adaptation.
@@ 3.2 Data-Driven Stability: Multi-Domain Experience Replay
  • General Domain Anchor ($\mathcal{B}_{gen}$): A gender-balanced subset of high-resource, standard Hindi samples (sourced from Kathbath) to preserve foundational phonetic robustness, which is essential for handling diverse patient demographics in a remote clinical setting.
@@ 3.3 Parameter-Driven Stability: Elastic Weight Consolidation (EWC)
We compute the importance of each LoRA parameter $\theta_{i}$ using the diagonal of the Fisher Information Matrix ($F$). To ensure robust optimization against high-variance gradients, we approximate importance using accumulated absolute gradients rather than squared gradients. This “Absolute Fisher” approach (hsu2023absolute) functions as a reliability optimization. We distinguish this from Memory Aware Synapses (MAS) (aljundi2018memory); while MAS assesses importance based on the sensitivity of the output function, Absolute Fisher derives importance from a linearized measure of the likelihood’s Fisher Information, preventing transient acoustic artifacts from drowning out meaningful weights:
@@ 4.1.2 Datasets
  • Gram Vaani (Clinical Helpline Proxy) (gram_vaani): This dataset consists of rural telephonic speech (originally 8kHz, upsampled to 16kHz) and serves as a proxy for audio encountered in frontline healthcare helplines. We evaluate on the official 3-hour Evaluation set (GV_Eval_3h). Note that prior work (patel2022gramvaani) reported 30.3% WER on the 5-hour Development set using full Conformer models. Our result (33.94%) reflects performance on the unseen Evaluation partition under strictly constrained on-device conditions (LoRA vs. full tuning). To strictly simulate a continual clinical learning scenario, we partition the 103 hours of training data into sequential segments, processing them one by one to mimic a live data stream while ensuring patient data residency.
@@ 5.2 Stability-Plasticity Pareto Efficiency
  • Data Anchors (ER): Replaying Kathbath samples provides an explicit gradient signal toward a generalized linguistic center. This was the most efficient stabilizer, as shown by the V3.1 Pareto trajectory.
  • Parameter Anchors (EWC): While EWC protects weights, it is “acoustically blind,” penalizing drift without reconciling new acoustic features. This is evidenced by V2.1 (ER) providing superior stabilization (+3.33% forgetting) compared to V4.5 (EWC, +3.58%) despite similar target performance.
  • The LoRA-Regularization Bottleneck: We observe a conflict in the Hybrid model (V5.1), where combining LoRA’s low-rank space with EWC’s parameter-level penalties overly restricts optimization paths, leading to higher target WER.
5.4 Convergence Dynamics
While the Naive baseline (V1.1) exhibits erratic loss curves, the ER and Hybrid paradigms demonstrate controlled, monotonic trends (Figure 2). This stability is a critical requirement for clinical deployment, ensuring the system does not regress after extended use. However, we observe a “soft floor” around 34% WER, which we characterize as the Acoustic Bottleneck of Clinical Telephony. The mismatch between 8kHz telephony and the 16kHz model expectation creates a ceiling for acoustic-only adaptation, suggesting that further gains may require specialized front-end processing or linguistic context.
@@ 5.5 The Necessity of Acoustic Adaptation
To verify that acoustic adaptation is a prerequisite for clinical deployment, we paired the baseline and adapted models with a domain-specific 4-gram LM (Table 2). The unadapted baseline+LM achieves only 34.96% WER, while V3.1+LM reaches 30.26% WER. This confirms that Acoustic Data Residency is the primary hurdle; solving for local acoustic mismatch is fundamental before linguistic correction can be effectively applied.
@@
5.6 Clinical Implications and Deployment Feasibility
The 17.1% relative improvement in WER demonstrated by our Multi-Domain ER strategy (V3.1) represents a significant bridge between "raw" and "usable" clinical ASR. In a clinical dictation setting, this reduction minimizes the manual correction burden on healthcare workers, whose time is a critical resource in rural India. By enabling the system to stabilize on local dialects and noisy line conditions without transmitting patient data to central servers, we provide a technical blueprint for scalable, privacy-compliant medical transcription. We note that prior work on the Gram Vaani dataset patel2022gramvaani achieved a lower WER (30.3%) using heavy-duty architectures (Conformer/TDNN) and external Language Models. However, our framework (33.94%) operates under strict on-device constraints without external LMs during training, reducing computational footprint while ensuring adherence to rigid data residency requirements. Furthermore, the ability to run these adaptations on mobile GPUs ensures that hospitals with limited digital infrastructure can still leverage state-of-the-art ASR foundational models.